# Nonmonotonic diffusion in sheared active suspensions of squirmers

### v1

@@ Abstract
We investigate how shear influences the dynamics of active particles in dilute to concentrated suspensions. Specifically, we focus on apolar active suspensions of squirmers, which are individually immotile but display activity-induced hydrodynamic diffusion. Under shear, the instantaneous particle velocities fluctuate more rapidly, similar to passive suspensions; however, surprisingly, the long-time diffusive dynamics can slow down and exhibit a nonmonotonic dependence on the shear rate. We show that this behavior arises from an interplay between the activity-induced persistent motion and shear-induced negative velocity autocorrelation, both of which are more pronounced at lower volume fractions. Simulations of self-propelled particles with tunable persistence support this explanation and further offer a simple mechanism to understand the observed coupling. Our results reveal a generic effect of shear on diffusion in active suspensions, elucidate how internal and external forcing interact, and provide new possibilities to modulate transport in active fluids.
@@
In this Letter, we investigate the interplay between shear and activity, and examine whether increasing shear rate enhances diffusion in active fluids. To this end, we use a recently modified Fast Stokesian Dynamics (FSD) method [30, 31] to perform large-scale hydrodynamic simulations of dilute to concentrated active suspensions ($\phi\in[10\%,40\%]$). Specifically, we simulate apolar active suspensions of “shakers,” particles that are active but not self-propelling, which collectively display activity-induced hydrodynamic diffusion [31]. The dynamics can be characterized by a diffusion coefficient, $D_{0}$, allowing for a Péclet number, Pe $\equiv\dot{\gamma}a^{2}/D_{0}$, to be defined to measure the relative effect of shear versus activity. Extensive simulations of thousands of shakers at different Pe and $\phi$ show that such suspensions always reach diffusive dynamics at long times; however, surprisingly, the effective diffusion coefficient, $D$, can drop below $D_{0}$ for a range of Pe when $\phi$ is small. For example, $D/D_{0}\approx 0.4$ when Pe $\approx 100$ and $\phi=10\%$. As Pe increases further, $D/D_{0}$ rises again, with shear eventually dominating, resulting in nonmonotonic diffusion. We analyze this behavior in detail and trace its physical origin to an activity-induced persistent motion and a shear-induced negative velocity autocorrelation, both of which are more pronounced at lower volume fractions. Additional simulations of self-propelled particles with tunable persistence support this interpretation and further provide a simple mechanism to understand the observed coupling. By contrast, the relative diffusivity of passive colloids with negligible persistence increases monotonically with Pe, as expected. Overall, our results demonstrate a general effect of shear on diffusion in active suspensions, elucidate how interactions between internal and external forcing influence these complex fluids, and suggest new ways to control mixing and mass transport in active fluids.
Figure 1: Nonmonotonic diffusion in sheared active suspensions. (a) Illustration of a typical simulation, showing 1024 shakers at 10% volume fraction under shear. (b) Mean square displacements in the $yz$-plane for pulling shakers or passive particles under different shear rates ($\phi=10\%$). The inset shows a few sample trajectories of shakers at $\dot{\gamma}=10^{-3}$ and 1. (c) Relative diffusion coefficient, $D/D_{0}$, against Pe for pulling or pushing shakers at different $\phi$.
@@
Although shakers are individually immotile, they swim collectively due to hydrodynamic interactions (HIs). To leading order, the HIs arise from $\bm{E}_{\text{s}}$ if $B_{2}\neq 0$: $B_{2}>0$ corresponds to pullers, while $B_{2}<0$ pushers. Therefore, the ratio $\beta\equiv B_{2}/B_{1}$ indicates the type of swimmer and the relative importance of HIs and self-propulsion. In this work, we mainly take $\beta\to\pm\infty$ to amplify the effect of HIs, but have also considered neutral squirmers with $\beta=0$ and additional rotational diffusion of $\bm{p}$ on each particle, effectively modeling active Brownian particles (ABPs, discussed later). Higher-order squirming modes and the azimuthal slip ($C_{1}$) are neglected, as is common in the literature [36, 37, 38, 31].
The dynamics of a suspension of squirmers is described by the active Stokesian Dynamics framework [35]. The governing equations can be written compactly as
@@
where the velocities ($\mathbf{U},\mathbf{U}_{\infty},\mathbf{U}_{\text{s}}$), strain rates ($\mathbf{E}_{\infty},\mathbf{E}_{\text{s}}$) and force ($\mathbf{F}_{\text{ext}}$) are all $N$-tuples, with elements defined at the particle centers, e.g., ${\mathbf{U}}\equiv(\bm{U}_{1}\ \bm{U}_{2}\ ...\ \bm{U}_{N})^{T}$, and ${\mathcal{R}}_{\text{FU}}$ and ${\mathcal{R}}_{\text{FE}}$ are the resistance tensors coupling the force and velocity moments of all $N$ particles. Note how individual particle activities couple to the large-scale flows through HIs. As illustrated in Fig. 1(a), we apply a shear in the $x$ direction, $\bm{U}_{\infty}(x,y,z)=(\dot{\gamma}y,0,0)$, and impose a short-range repulsive force $\bm{F}_{\text{ext}}$ between the particles to model excluded-volume interactions. Eq. (3) is then solved using FSD [30] for $N=1024$ at different Pe and $\phi$; see the supplemental material (SM)  [^1] See Supplemental Material at [URL will be inserted by publisher] for a detailed description of our simulation method, setup and verification, which includes Refs. [50, 51, 52, 53]. for further details.
@@
Apolar active suspensions—Recently, we showed that suspensions of shakers, pulling or pushing, could undergo activity-induced diffusion without any external force [31]. The effective diffusion coefficient was found to be $D_{0}=(a^{2}/\tau_{a})g(\phi)$, where $\tau_{a}\equiv a/|B_{2}|$ is the activity time scale, and $g(\phi)\sim\phi^{1/5}$ for $\phi\lessapprox 10\%$ and reduces sharply with $\phi$ beyond. Furthermore, the short-time dynamics was ballistic and could last over tens of $\tau_{a}$; c.f. Fig. 1(b). This persistent motion, therein due to HIs between the active particles, is the main reason that active fluids display an enhanced diffusion [3, 6, 31].
@@
Figure 2: Short-time dynamics and phase diagram. (a) Root-mean-square particle speeds under shear, normalized by their values without shear at all Pe and $\phi$. (b) Normalized velocity autocorrelation for passive particles at different $\phi$ (markers denote $\Phi_{v}(\dot{\gamma}\tau_{d})=-10^{-2}$) and for pulling shakers under different Pe at $\phi=10\%$ (inset). (c) Phase diagram of all simulation results for shakers, divided by Pe $=1$ and 100 into three diffusive regimes with different scalings.
@@
We extract the diffusion coefficient, $D$, by fitting the long-time MSD according to $\langle\Delta y^{2}(t)+\Delta z^{2}(t)\rangle=4D(t-t_{0})$, where $t_{0}$ is generally nonzero due to nondiffusive dynamics at short times. Fig. 1(c) shows the relative diffusivity, $D/D_{0}$, across four decades of Pe. Clearly, $D/D_{0}$ is nonmonotonic at $\phi=10\%$, with a minimum of about 0.4 reached at Pe $\approx 100$, consistent with the observation in Fig. 1(b). At higher concentrations, the nonmonotonic behavior gradually disappears, and the scaling between $D/D_{0}$ and Pe depends on $\phi$. However, in all cases, we observe $D/D_{0}\approx 1$ for Pe $\lessapprox 1$ and $D/D_{0}\propto$ Pe for Pe $\gtrapprox 100$. The same scalings apply to both pullers and pushers, as is the case without shear [31], implying similar but still irreversible macroscopic dynamics under time reversal.
To explain the nonmonotonic diffusion, we examine the short-time dynamics characterized by the instantaneous particle velocities, projected onto the $yz$-plane as in the MSD calculation. Fig. 2(a) shows the root-mean-square particle speeds under shear, $U_{rms}$, normalized by their values without shear, $U_{rms,0}$, for pullers and pushers at different $\phi$. Remarkably, all data collapse onto a single curve, showing universal scalings with $U_{rms}/U_{rms,0}\approx 1$ for Pe $\lessapprox 1$ and $U_{rms}/U_{rms,0}\propto$ Pe for Pe $\gtrapprox 10$. The absence of any $\phi$-dependence suggests that the relative change of the short-time dynamics is controlled by shear alone, despite the underlying activity-induced dynamics, $U_{rms,0}$ and $D_{0}$, depend nonmonotonically on $\phi$ [31].
[…]
We now explain the observed diffusion. Fig. 2(c) shows the phase diagram of all shaker simulations controlled by two parameters, $\phi$ and $\dot{\gamma}\tau_{a}=g(\phi)$Pe. As discussed, there are three diffusive regimes separated by Pe $\approx 1$ and 100, with different scalings between $D/D_{0}$ and Pe in each. Specifically, Pe $<1$ corresponds to $a^{2}/D_{0}<\dot{\gamma}^{-1}$, i.e., the activity-induced diffusion occurs faster than the shear-induced, thus we observe the former, $D=D_{0}$. When Pe $>1$, both activity and shear are at play, resulting in $\phi$-dependent scalings, including a reduced diffusion at low $\phi$ due to the shear-induced velocity anticorrelation. However, at sufficiently large Pe, the dynamics must be controlled by shear, $D\propto\dot{\gamma}a^{2}$, leading to $D/D_{0}\propto$ Pe. Here, the threshold Pe corresponds to $\tau_{d}=\tau_{a}$, where $\tau_{d}$ is the velocity decorrelation time, i.e., the time $\Phi_{v}(t)$ returns to zero from below, c.f. Fig. 2(b). When $\tau_{d}<\tau_{a}$, diffusion occurs due to shear because even the shortest activity time scale ($\tau_{a}$) is longer than the diffusive time of shear ($\tau_{d}$). Indeed, Fig. 2(c) shows that $\tau_{a}=\tau_{d}$ approximates Pe $=100$, which confirms our analysis.
ABPs—One may wonder whether the nonmonotonic diffusion observed here is a general feature of fluids with persistent internal motion. After all, if it is observable in apolar active suspensions, the effect may be stronger for polar particles with a larger persistence. To test this hypothesis, we perform additional simulations of ABPs, a widely studied model [41] for self-propelled particles, e.g., self-phoretic colloids [42]. Specifically, we simulate neutral squirmers with $B_{1}>0$ and $B_{2}=0$, and impose a rotational diffusion, $d_{r}$, to effect angular noise in the particle dynamics, $\langle\bm{\Omega}_{\text{s}}(t)\cdot\bm{\Omega}_{\text{s}}(t^{\prime})\rangle=2d_{r}\delta(t-t^{\prime})$, where $\delta(t)$ is the Dirac delta. Translational noise is neglected as done typically [43]. Without shear, our ABPs self-propel with a constant speed, $U_{\text{s}}=\frac{2}{3}B_{1}$, unless they collide. The underlying diffusivity is $D_{0}=\frac{1}{3}U_{\text{s}}^{2}\tau_{r}$, where $\tau_{r}=\frac{1}{2}d_{r}^{-1}$, in the dilute regime (see SM [39]), and the persistence length, $\ell_{p}\equiv U_{\text{s}}\tau_{r}$, is tunable.
Figure 3: Diffusivity of ABPs with different persistence length ($\ell_{p}$) at $\phi=10\%$. The main plot shows the relative diffusivity in the $y$ direction; the inset in $z$.
@@
Fig. 3 shows the relative diffusivities, $D_{yy}/D_{0}$ and $D_{zz}/D_{0}$, in $y$ and $z$ directions, separately, for ABPs with different $\ell_{p}$ at $\phi=10\%$ under shear. The diffusivities are extracted in the same way as for shakers. For $D_{yy}/D_{0}$, we observe the expected effect of $\ell_{p}$, i.e., a larger persistence leads to a larger drop in diffusivity, which can be more than 90% if $\ell_{p}/a>3$ (as a reference, $\ell_{p}/a\approx 1.4$ for shakers at $\phi=10\%$ [31]), and the nonmonotonic effect disappears when $\ell_{p}/a\ll 1$. Conversely, $D_{zz}/D_{0}$ always increases with Pe, and the data at different $\ell_{p}$ collapse, scaling as SID. We observe these contrasting behaviors because ABPs hardly interact. As shear rotates all particles around the $z$ axis with a constant speed $\Omega_{\infty}=\dot{\gamma}/2$, ABPs oscillate in the $xy$-plane but remain persistent along $z$ (manifest in the MSDs, see SM [39]). Consequently, $D_{zz}/D_{0}$ can decouple from $D_{yy}/D_{0}$ over a large range of Pe, leading to highly anisotropic diffusion.
The above results depict a simple physical picture of the interplay between shear and activity. When shearing a fluid with an underlying persistence, arising from either activity-induced HIs (as in the case of shakers), self-propulsion (ABPs), or a combination of the two (squirmers in general), the rotational effect of the shear reduces the persistent motion in the shear plane. In the dilute regime where SID is weak, reduced persistence leads to reduced diffusion, as exemplified by a single ABP under shear [44, 45]. However, as the shear rate increases, particles collide more often due to stronger advection, eventually resulting in increased diffusion. The exact behavior depends on the propulsion and interaction mechanisms; nevertheless, the competition between external rotation, internal persistence, and SID determines the diffusive dynamics of active fluids in general.
Figure 4: Diffusivity of passive colloidal suspensions at different $\phi$. The inset shows the MSDs at $\phi=10\%$ under different Pe, where $\tau_{B}\equiv a^{2}/\big{(}k_{\text{B}}T\mu(0)\big{)}$.
Passive colloidal suspensions—The dynamics of ABPs suggests that a persistent internal motion is necessary for observing a nonmonotonic diffusion under shear. To prove this, we consider the contraposition, i.e., passive colloids without any persistence. The underlying diffusivity is $D_{0}=k_{\text{B}}T\mu(\phi)$, where $k_{\text{B}}$ is the Boltzmann’s constant, $T$ the absolute temperature, and $\mu(\phi)$ a decreasing function of $\phi$ [46]. Fig. 4 shows the relative diffusivity, $D/D_{0}$, against Pe for $\phi$ from 10% to 40%, as well as the MSDs under different Pe at $\phi=10\%$. Clearly, shear tends to enhance the diffusion at all $\phi$, despite the subdiffusive dynamics appearing transiently at higher Pe (see Fig. 4 inset and SM [39]). These observations thus verify our hypothesis, highlighting a fundamental difference between active and passive fluids under shear.
Concluding remarks—We have presented a numerical study on the dynamics of sheared particle suspensions using rigorous hydrodynamic simulations. The central finding is that, while shear always enhances the dynamics of passive particles, it can suppress the diffusion in active suspensions by reducing the persistent motion of active particles. This interplay between shear and activity is more pronounced in dilute suspensions, where the activity-induced diffusivity is larger than the shear-induced, and it does not depend on the origin of the persistent motion, being either an intrinsic property of the particles or an emergent phenomenon due to HIs. Therefore, we expect the observed coupling to be a general feature of fluids endowed with an underlying microstructure and internal persistence. Finally, there is currently a growing interest in the relation between the dynamics and rheology of active fluids, which has great practical implications but is still not well understood [45, 47, 48, 49]. Our study may shed some light on this problem.
@@
Acknowledgments—We are grateful to B. Chen for code development, and thank E. Kanso and T. Zhou for discussions. Z.G. acknowledges support from the Swedish Research Council under Grant No. 2021-06669VR.

### v2

@@ Abstract
We investigate how shear influences the dynamics of active particles in dilute to concentrated suspensions. Using apolar active suspensions of squirmers as model systems, we show how their long-time diffusive dynamics can surprisingly slow down and vary nonmonotonically with the shear rate arising from an interplay between the activity-induced persistent motion and shear-induced reorientation and diffusion. Further simulations of self-propelled particles with tunable persistence exhibit richer dynamics and confirm the observed coupling, suggesting that nonmonotonic diffusion may be a general feature of fluids endowed with an underlying microstructure and large persistence. Our results reveal a nonlinear effect of shear on diffusion in active suspensions, elucidate how internal and external forcing interact, and provide new possibilities to modulate transport in active fluids.
@@
In this Letter, we investigate the interplay between shear and activity, and examine whether increasing shear rate enhances diffusion in active fluids. To this end, we use a recently modified Fast Stokesian Dynamics (FSD) method [30, 31] to perform large-scale hydrodynamic simulations of dilute to concentrated active suspensions ($\phi\in[10\%,40\%]$). Specifically, we first consider a simple model – apolar active suspensions of “shakers,” particles that are active but not self-propelling, which collectively display activity-induced hydrodynamic diffusion [31]. The dynamics can be characterized by a diffusion coefficient, $D_{0}$, allowing for a Péclet number, Pe $\equiv\dot{\gamma}a^{2}/D_{0}$, to be defined to measure the relative effect of shear versus activity. Extensive simulations of thousands of shakers at different Pe and $\phi$ show that such suspensions always reach diffusive dynamics at long times; however, surprisingly, the effective diffusion coefficient, $D$, can drop below $D_{0}$ for a range of Pe when $\phi$ is small. For example, $D/D_{0}\approx 0.4$ when Pe $\approx 100$ and $\phi=10\%$. As Pe increases further, $D/D_{0}$ rises again, with shear eventually dominating, resulting in nonmonotonic diffusion. We analyze this behavior in detail and trace its physical origin to an activity-induced persistent motion and a shear-induced negative velocity autocorrelation, both of which are more pronounced at lower volume fractions. To confirm and generalize our understanding, we then consider self-propelled particles with tunable persistence, which display nonmonotonic diffusion even in concentrated suspensions, provided that the persistence is large enough. Overall, our results demonstrate a nonlinear effect of shear on diffusion in active suspensions, elucidate how interactions between internal and external forcing influence these complex fluids, and suggest new ways to control mixing and mass transport in active fluids.
Figure 1: Nonmonotonic diffusion in sheared active suspensions. (a) Illustration of 1024 particles at 10% volume fraction ($\phi$) under shear. (b) Mean square displacements in the $yz$-plane for pulling shakers ($\phi=10\%$) at different Péclet numbers, Pe. The inset shows a few sample trajectories at Pe =0=0 and 77.5. (c) Relative diffusion coefficient, $D/D_{0}$, against Pe for pulling or pushing shakers at different $\phi$. Slopes are theoretical estimates (see text).
@@
Although shakers are individually immotile, they swim collectively due to hydrodynamic interactions (HIs). To leading order, the HIs arise from $\bm{E}_{\text{s}}$ if $B_{2}\neq 0$: $B_{2}>0$ corresponds to pullers, while $B_{2}<0$ pushers. Therefore, the ratio $\beta\equiv B_{2}/B_{1}$ indicates the type of swimmer and the relative importance of HIs and self-propulsion. In this work, we first take $\beta\to\pm\infty$ to amplify the effect of HIs, then consider neutral squirmers with $\beta=0$ and additional rotational diffusion of $\bm{p}$ on each particle, effectively modeling active Brownian particles (ABPs). Higher-order squirming modes and the azimuthal slip ($C_{1}$) are neglected, as is common in the literature [36, 37, 38].
The dynamics of a suspension of squirmers is described by the active Stokesian Dynamics framework [35]. For athermal particles, the governing equations can be written compactly as
@@
where the velocities ($\mathbf{U},\mathbf{U}_{\infty},\mathbf{U}_{\text{s}}$), strain rates ($\mathbf{E}_{\infty},\mathbf{E}_{\text{s}}$) and force ($\mathbf{F}_{\text{ext}}$) are all $N$-tuples, with elements defined at the particle centers, e.g., ${\mathbf{U}}\equiv(\bm{U}_{1}\ \bm{U}_{2}\ ...\ \bm{U}_{N})^{T}$, and ${\mathcal{R}}_{\text{FU}}$ and ${\mathcal{R}}_{\text{FE}}$ are the resistance tensors coupling the force and velocity moments of all $N$ particles. Note how individual particle activities couple to the large-scale flows through HIs. As illustrated in Fig. 1(a), we apply a shear in the $x$ direction, $\bm{U}_{\infty}(x,y,z)=(\dot{\gamma}y,0,0)$, and impose a short-range repulsive force $\bm{F}_{\text{ext}}$ between the particles to model excluded-volume interactions; see the Supplemental Material (SM)  [^1] See Supplemental Material at [url] for a detailed description of our simulation method, setup and verification, which also includes Refs. [52, 53, 54, 55, 56, 57, 58]. for further details. Eq. (3) is then solved using FSD [30] for $N=1024$ at different Pe and $\phi$.
@@
Shakers—Recently, we showed that apolar active suspensions of shakers, pulling or pushing, could undergo activity-induced diffusion without any external flow [31]. The effective diffusion coefficient was found to be $D_{0}=(a^{2}/\tau_{a})g(\phi)$, where $\tau_{a}\equiv a/|B_{2}|$ is the activity time scale, and $g(\phi)\sim\phi^{1/5}$ for $\phi\lessapprox 10\%$ and reduces sharply with $\phi$ beyond. Furthermore, the short-time dynamics was ballistic and could last over tens of $\tau_{a}$; see the Pe =0=0 case in Fig. 1(b). This persistent motion, therein due to HIs between the active particles, is the main reason that active fluids display an enhanced diffusion [3, 6, 31].
@@
Figure 2: Short-time dynamics and phase diagram. (a) Root-mean-square particle speeds under shear, normalized by their values without shear at all Pe and $\phi$. (b) Normalized velocity autocorrelation for passive particles at different $\phi$ (markers denote $\Phi_{v}(\dot{\gamma}\tau_{d})=-10^{-2}$) and for pulling shakers under different Pe at $\phi=10\%$ (inset). (c) Phase diagram of all simulation results for shakers, divided by Pe =1=1 and 100 into three diffusive regimes with different scalings.
Under shear, the particle dynamics is anisotropic and not diffusive in the flow direction ($x$); thus, we analyze the mean square displacement (MSD) in the $yz$-plane, $\langle\Delta y^{2}(t)+\Delta z^{2}(t)\rangle$, where $\langle\cdot\rangle$ averages over all particles and time. Fig. 1(b) shows the MSDs for pulling shakers under different Pe at $\phi=10\%$ (see SM for results of pushers and at other $\phi$ [39]). We observe that the short-time MSD increases with Pe, similar to the behavior of passive suspensions [23]; however, surprisingly, the long-time dynamics can be suppressed by shear, with sustained subdiffusive motion, resulting in smaller MSDs at larger Pe. The reduced diffusion is also evident from direct visualization of individual particle trajectories. As shown in the inset in Fig. 1(b), increasing Pe could lead to more compact and less smooth trajectories at long times.
We extract the diffusion coefficient, $D$, by fitting the long-time MSD according to $\langle\Delta y^{2}(t)+\Delta z^{2}(t)\rangle=4D(t-t_{0})$, where $t_{0}$ is generally nonzero due to nondiffusive dynamics at short times. Fig. 1(c) shows the relative diffusivity, $D/D_{0}$, across four decades of Pe. Clearly, $D/D_{0}$ is nonmonotonic at $\phi=10\%$, with a minimum of about 0.4 reached at Pe $\approx 100$, consistent with the observation in Fig. 1(b). At higher concentrations, the nonmonotonic behavior gradually disappears, and the scaling between $D/D_{0}$ and Pe depends on $\phi$ (see analysis later). However, in all cases, we observe $D/D_{0}\approx 1$ for Pe $\lessapprox 1$ and $D/D_{0}\propto$ Pe for Pe $\gtrapprox 100$. The same scalings apply to both pullers and pushers, as is the $\phi$-dependence of $D_{0}$ without shear (c.f. Fig. 4) [31], implying similar but still irreversible macroscopic dynamics under time reversal.
To explain the nonmonotonic diffusion, we examine the short-time dynamics characterized by the instantaneous particle velocities, projected onto the $yz$-plane as in the MSD calculation. Fig. 2(a) shows the root-mean-square particle speeds under shear, $U_{rms}$, normalized by their values without shear, $U_{rms,0}$, for pullers and pushers at different $\phi$. Remarkably, all data collapse onto a single curve, showing universal scalings with $U_{rms}/U_{rms,0}\approx 1$ for Pe $\lessapprox 1$ and $U_{rms}/U_{rms,0}\propto$ Pe for Pe $\gtrapprox 10$. The absence of any $\phi$-dependence suggests that the relative change of the short-time dynamics is controlled by shear alone, despite the underlying activity-induced dynamics, $U_{rms,0}$ and $D_{0}$, depending nonmonotonically on $\phi$ [31].
The instantaneous velocities decorrelate over time and can be quantified by a velocity autocorrelation function, $\Phi_{v}(t)\equiv\langle\bm{U}(\tau)\cdot\bm{U}(\tau+t)\rangle/U_{rms}^{2}$, where $\langle\cdot\rangle$ averages over all particles and time. Fig. 2(b) shows $\Phi_{v}(t)$ for pulling shakers under different Pe at $\phi=10\%$ and passive particles (Pe $\to\infty$ as activity is turned off) at different $\phi$. We observe faster decorrelation as Pe or $\phi$ increases and even a negative $\Phi_{v}(t)$ at sufficiently high Pe. Such anticorrelation has been noted in earlier simulations [26, 23] and results from either hydrodynamic or excluded-volume interactions. Under shear, pairwise HIs render transient particle displacements that necessarily become zero at equal time after the encounter [40]; thus, the velocities change sign and $\Phi_{v}(t)<0$. This effect is weakened at higher $\phi$ due to many-body HIs and more frequent collisions; however, particles tend to oscillate as they collide (see supplemental videos), again making $\Phi_{v}(t)<0$. The time scale of these shear-induced interactions is proportional to $\dot{\gamma}^{-1}$ and reduces with $\phi$, c.f. Fig. 2(b). The negative velocity autocorrelation is the reason for the subdiffusive motion observed in Fig. 1(b).
We now explain the observed diffusion. Fig. 2(c) shows the phase diagram of all shaker simulations controlled by two parameters, $\phi$ and $\dot{\gamma}\tau_{a}=g(\phi)$Pe. As discussed, there are three diffusive regimes separated by Pe $\approx 1$ and 100, with different scalings between $D/D_{0}$ and Pe in each. Specifically, Pe <1<1 corresponds to $a^{2}/D_{0}<\dot{\gamma}^{-1}$, i.e., the activity-induced diffusion occurs faster than the shear-induced, thus we observe the former, $D=D_{0}$. When Pe >1>1, both activity and shear are at play, resulting in $\phi$-dependent scalings, including a reduced diffusion at low $\phi$ due to the shear-induced velocity anticorrelation. However, at sufficiently large Pe, the dynamics must be controlled by shear, $D=D_{s}\propto\dot{\gamma}a^{2}$, leading to $D/D_{0}\propto$ Pe. Here, the threshold Pe corresponds to $\tau_{d}=\tau_{a}$, where $\dot{\gamma}\tau_{d}$ is a $\phi$-dependent velocity decorrelation time, i.e., the time $\Phi_{v}(t)$ returns to zero from below, c.f. Fig. 2(b). When $\tau_{d}<\tau_{a}$, diffusion occurs due to shear because even the shortest activity time scale ($\tau_{a}$) is longer than the diffusive time of shear ($\tau_{d}$). Indeed, Fig. 2(c) shows that $\tau_{a}=\tau_{d}$ approximates Pe =100=100. Furthermore, we can also estimate the scaling between $D/D_{0}$ and Pe in the intermediate Pe range. Suppose $D/D_{0}\sim$ Peν, and let $D=D_{0}$ (or $D_{s}$) at Pe == Pe0 (or Pes); then, $\nu=\log(D_{s}/D_{0})/\log(\textrm{Pe}_{s}/\textrm{Pe}_{0})$. Using $_{0}=1, Pe$$_{s}=100$, and $D_{s}/D_{0}=0.4$ (or 4.0), we obtain $\nu=-0.2$ (or 0.3) at $\phi=0.1$ (or 0.4). Fig. 1(c) verifies our estimates.
ABPs—One may wonder whether the nonmonotonic diffusion observed here is a general feature of fluids with persistent internal motion. After all, if it is observable in apolar active suspensions, the effect may be stronger for polar particles with a larger persistence. To test this hypothesis, we perform additional simulations of ABPs, a widely studied model [41] for self-propelled particles, e.g., self-phoretic colloids [42]. Specifically, we simulate neutral squirmers with $B_{1}>0$ and $B_{2}=0$, and impose a uniform rotational diffusion, $d_{r}$, to effect angular noise in the particle dynamics, $\langle\bm{\Omega}_{\text{s}}(t)\cdot\bm{\Omega}_{\text{s}}(t^{\prime})\rangle=2d_{r}\delta(t-t^{\prime})$, where $\delta(t)$ is the Dirac delta. Translational noise is neglected as done typically [43]. Without shear, our ABPs self-propel with a constant speed, $U_{s}=\frac{2}{3}B_{1}$, unless they collide. Over time, collision and orientation relaxation render the motion diffusive, with $D_{0,\textrm{ABPs}}\to\frac{1}{3}U_{s}^{2}\tau_{r}$ as $\phi\to 0$, where $\tau_{r}\equiv\frac{1}{2}d_{r}^{-1}$. The persistence length, $\ell_{p}\equiv 3D_{0,\textrm{ABPs}}/U_{rms}$, is tunable and capped at $\ell_{p,0}\equiv U_{s}\tau_{r}$ as $\phi\to 0$ [39].
@@
Fig. 3 shows the separate diffusivities in $y$ and $z$ directions, $D_{yy}$ and $D_{zz}$, relative to $D_{0,\textrm{ABPs}}$, for different $\ell_{p}\approx\ell_{p,0}$ [39] at $\phi=10\%$ under shear. The diffusivities are extracted in the same way as for shakers. For $D_{yy}$, we observe the expected effect of $\ell_{p}$, i.e., a larger persistence leads to a larger drop in diffusivity, which can be more than 90% if $\ell_{p}/a>3$ (as a reference, $\ell_{p}/a\sim O(1)$ for shakers at $\phi=10\%$ [31]). Conversely, the nonmonotonic effect disappears when $\ell_{p}/a\ll 1$, and we have verified that the diffusivity of passive colloids ($\ell_{p}=0$) strictly increases with shear [39]. The Pe at which motility ceases (or shear begins) to dominate, Pe0 (or Pes), can be estimated similarly as in the case of shakers: motility dominates when $\max(a^{2}/D_{0,\textrm{ABPs}},\ell_{p}^{2}/D_{0,\textrm{ABPs}})<\dot{\gamma}^{-1}$, or Pe $\min(1,(a/\ell_{p})^{2})\equiv$ Pe0, while shear dominates when $\tau_{d}<\min(\tau_{s},\tau_{r})$, or Pe $\dot{\gamma}_{s}a^{2}/D_{0,\textrm{ABPs}}\equiv$ Pes, where $\tau_{s}\equiv a/U_{s}$ and $\dot{\gamma}_{s}\equiv\dot{\gamma}\tau_{d}/\min(\tau_{s},\tau_{r})$. Here, a major difference between ABPs and shakers is that $D_{zz}/D_{0,\textrm{ABPs}}$ always increases with Pe, and the data at different $\ell_{p}$ collapse, scaling as SID. The reason for the contrasting behaviors is that ABPs hardly interact. As shear rotates all particles around the $z$ axis with a constant speed $\Omega_{\infty}=\dot{\gamma}/2$, ABPs oscillate in the $xy$-plane but remain persistent along $z$ (manifest in the MSDs [39]). Consequently, $D_{zz}/D_{0,\textrm{ABPs}}$ can decouple from $D_{yy}/D_{0,\textrm{ABPs}}$ over a large range of Pe, leading to highly anisotropic diffusion.
Figure 4: Comparison of the different intrinsic and shear-induced diffusivities (in the same unit, $a^{2}/\tau_{a}$). The main plot compares the $D_{0}$ for pulling or pushing shakers with $D_{s}$ for passive particles at $\dot{\gamma}_{s}(\phi=0.2)$. The inset compares $D_{0,\textrm{ABPs}}$ at different $\ell_{p,0}$ (same color legend as in Fig. 3) with $D_{s}$ at $\dot{\gamma}_{s}(\phi=0.4,\ell_{p,0}/a=10)$.
Discussions—The results of our simulations depict a general physical picture of the interplay between shear and activity. When shearing a fluid with an underlying persistence, arising from either activity-induced HIs (as in the case of shakers), self-propulsion (ABPs), or a combination of the two (squirmers in general), the rotational effect of shear reduces the persistent motion in the shear plane. For a single ABP without any SID (as $\phi\to 0$), this leads to a monotonically reducing $D_{yy}/D_{0,\textrm{ABPs}}$ with $\dot{\gamma}\tau_{r}$ [44, 45]. At finite $\phi$, both shear-induced rotation and collision are at play, and the dynamics depends on the competition of the intrinsic diffusion and SID. If $D_{s}(\dot{\gamma}_{s},\phi)<D_{0}(\tau_{a},\phi)$, where $\dot{\gamma}_{s}(\phi)$ is the shear rate above which SID dominates, we would expect a reduced diffusion as Pe increases from Pe0 to Pes. Conversely, if $D_{s}(\dot{\gamma}_{s},\phi)>D_{0}(\tau_{a},\phi)$, an enhanced diffusion is expected. Fig. 4 shows that $D_{0}$ of both pulling and pushing shakers cross $D_{s}$ at $\phi\approx 20\%$, which is why $D/D_{0}$ is nearly independent of Pe up to Pes at $\phi=20\%$; c.f. Fig. 1(c). Furthermore, the inset in Fig. 4 suggests that diffusion could be suppressed even in concentrated suspensions, since $D_{s}(\dot{\gamma}_{s},\phi)<D_{0,\textrm{ABPs}}(\tau_{s},\tau_{r},\phi)$ at sufficiently large $\ell_{p}$, which we indeed observe at $\phi=40\%$ and $\ell_{p,0}/a=10$ [39]. Therefore, we expect nonmonotonic diffusion to be a general feature when shearing fluids endowed with an underlying microstructure and large internal persistence. Finally, there is currently a growing interest in the relation of the dynamics, structure, and rheology of active fluids, which has great practical implications but is still not well understood [46, 45, 47, 48, 49, 50]. Our study may shed some light on this problem.
@@
Acknowledgments—We are grateful to Boyuan Chen for help in the code development, and thank Eva Kanso, Tingtao Zhou, and the anonymous referees for useful comments. Z.G. acknowledges support from the Swedish Research Council under Grant No. 2021-06669VR.
Data availability—The source code to generate all data reported in this article is publically available  [^2] https://github.com/GeZhouyang/FSD.